OpenStudy (anonymous):

Choose the equation that contains the point (8,0) and 0,2) Y=-1/4+2x. y=-1/4x+2. Y=1/4x-2

OpenStudy (vijay):

y=-1/4x+2.

OpenStudy (aroub):

2-0/0-8 2/-8 1/-4 y=mx+b , m= -1/4 , ( 0,2) 2=-1/4(0)+b 2=0+b 2=b so the equation is y=-1/4x +2